# Real $K$-Theory for $C^{*}$-Algebras: just the facts October 22, 2025

### v2

@@ 6. The Homogeneous Spaces $Sp_{n}/(Sp_{k}\times Sp_{n-k})$
We will see in the next section that $\pi_{j}(Sp_{n})=0$ for $j\leq 2$. This implies that in the sequence (*) that two out of every three groups is zero, and hence
@@ 8. The Homogeneous Spaces $Sp_{n}/U_{n}$ and $Sp_{n}/SU_{n}$
The spaces $Sp_{n}/U_{n}$ are path-connected. Sequence (**) gives an exact sequence
@@
For $Sp_{n}/SU_{n}$ sequence (**) gives an exact sequence
@@ 9. The Homogeneous Spaces $U_{n}/O_{n}$, $SU_{n}/SO_{n}$ and $U_{n}/SO_{n}$
First consider $U_{n}/O_{n}$. As $O_{n}$ is not path-connected we use sequence (*) and obtain
@@
shows that the map $\alpha:\pi_{1}(O_{2})\to\pi_{1}(U_{2})$ is an isomorphism. Thus the sequence (*) reduces down to
@@
$\pi_{1}(U_{n}/O_{n})\cong{\mathbb{Z}}_{2}\qquad and\qquad\pi_{2}(U_{n}/O_{n})=0.$
@@
The map $\alpha\equiv 0$ and so
$\pi_{2}(U_{n}/O_{n})={\mathbb{Z}}_{2}.$
We are left with
$0\to{\mathbb{Z}}\cong\pi_{1}(U_{n})\overset{\beta}{\longrightarrow}\pi_{1}(U_{n}/O_{n})\to{\mathbb{Z}}_{2}\to 0.$
This implies algebraically that $\pi_{1}(U_{n}/O_{n})$ is either $\mathbb{Z}$} with $\beta$ being multiplication by $2$,or ${\mathbb{Z}}\oplus{\mathbb{Z}}_{2}$ with $\beta$ being isomorphic inclusion onto the first factor. We will settle this point in a moment.
@@
Turning to sequence (**) we obtain
@@
Finally, consider $U_{n}/SO_{n}$. The space $U_{n}$ is path-connected. If $n=2$ then sequence (**) gives us
@@
$\pi_{1}(U_{n}/SO_{n})\cong{\mathbb{Z}}\qquad and\qquad\pi_{2}(U_{n}/SO_{n})\cong{\mathbb{Z}}_{2}$
@@
The map $\gamma$ is onto and hence $\pi_{0}(O_{n}/(O_{k}\times O_{n-k}))=0$. A similar and easier argument works for $SO$ and hence
$\pi_{0}(O_{n}/(O_{k}\times O_{n-k}))=0\qquad and\qquad\pi_{0}(SO_{n}/(SO_{k}\times SO_{n-k}))=0$
@@
    $\pi_{1}(O_{2}/(O_{1}\times O_{1}))\cong{\mathbb{Z}}\qquad and\qquad\pi_{2}(O_{2}/(O_{1}\times O_{1})=0.$
@@
    $\pi_{1}(SO_{2}/(SO_{1}\times SO_{1}))\cong{\mathbb{Z}}\qquad and\qquad\pi_{2}(SO_{2}/(SO_{1}\times SO_{1})=0.$
@@
    with $p$ onto, and this implies that
    $\pi_{1}(RP^{3}/S^{1})=0\qquad and\qquad\pi_{2}(RP^{3}/S^{1})={\mathbb{Z}}$
@@
    $\pi_{1}(O_{3}/O_{2})=0\qquad and\qquad\pi_{2}(O_{3}/O_{2})={\mathbb{Z}}$
@@
    $\pi_{1}(SO_{3}/SO_{2})=0\qquad and\qquad\pi_{2}(SO_{3}/SO_{2})={\mathbb{Z}}$
@@
    $\pi_{1}(O_{3}/(O_{1}\times O_{2})={\mathbb{Z}}_{2}\qquad and\qquad\pi_{2}(O_{3}/(O_{1}\times O_{2})\cong{\mathbb{Z}}.$
@@
    $\pi_{1}(SO_{3}/(SO_{1}\times SO_{2})=0\qquad and\qquad\pi_{2}(SO_{3}/(SO_{1}\times SO_{2})\cong{\mathbb{Z}}.$
@@
    Applying the sequence (*) to $O_{4}/(O_{2}\times O_{2})$ gives us
@@
    $0\to\pi_{2}(O_{4}/(O_{2}\times O_{2}))\to{\mathbb{Z}}^{2}\overset{onto}{\longrightarrow}{\mathbb{Z}}_{2}\overset{p}{\longrightarrow}\pi_{1}(O_{4}/(O_{2}\times O_{2}))\to{\mathbb{Z}}_{2}\oplus{\mathbb{Z}}_{2}\to{\mathbb{Z}}_{2}\to 0$
@@
    $0\to\pi_{2}(O_{4}/(O_{2}\times O_{2}))\to{\mathbb{Z}}^{2}\overset{onto}{\longrightarrow}{\mathbb{Z}}_{2}\to 0$
@@
    $\pi_{1}(O_{4}/(O_{2}\times O_{2}))\cong{\mathbb{Z}}_{2}\qquad and\qquad\pi_{2}(O_{4}/(O_{2}\times O_{2}))\cong{\mathbb{Z}}^{2}$
@@
    $0\to\pi_{2}(SO_{4}/(SO_{2}\times SO_{2}))\longrightarrow\pi_{1}(SO_{2}\times SO_{2})\overset{onto}{\longrightarrow}\pi_{1}(SO_{4})\longrightarrow\pi_{1}(SO_{4}/(SO_{2}\times SO_{2}))\to 0$
@@
    $0\to\pi_{2}(SO_{4}/(SO_{2}\times SO_{2}))\longrightarrow{\mathbb{Z}}^{2}\overset{onto}{\longrightarrow}{\mathbb{Z}}_{2}\overset{p}{\longrightarrow}\pi_{1}(SO_{4}/(SO_{2}\times SO_{2}))\to 0$
@@
    $\pi_{1}(SO_{4}/(SO_{2}\times SO_{2}))\cong 0\qquad and\qquad\pi_{2}(SO_{4}/(SO_{2}\times SO_{2}))\cong{\mathbb{Z}}^{2}$
@@
    The map $\bar{\gamma}$ is onto, and hence $\zeta$ is the zero map. The map $\gamma$ is also onto. Thus
    $\pi_{1}(O_{n}/(O_{k}\times O_{n-k}))\cong{\mathbb{Z}}_{2}\qquad and\qquad\pi_{2}(O_{n}/(O_{k}\times O_{n-k}))\cong{\mathbb{Z}}_{2}.$
@@
    The case of $SO_{n}/(SO_{k}\times SO_{n-k})$ is much simpler. The space $SO(n)$ is path-connected and sequence (**) gives us
@@
    $0\to\pi_{2}(SO_{n}/(SO_{k}\times SO_{n-k}))\longrightarrow{\mathbb{Z}}_{2}\oplus{\mathbb{Z}}_{2}\overset{onto}{\longrightarrow}{\mathbb{Z}}_{2}\to\pi_{1}(SO_{n}/(SO_{k}\times SO_{n-k}))\to 0.$
@@
    $\pi_{1}(SO_{n}/(SO_{k}\times SO_{n-k}))=0\qquad and\qquad\pi_{2}(SO_{n}/(SO_{k}\times SO_{n-k}))={\mathbb{Z}}_{2}.$
@@ 13. The Basics: Real Vector Bundles
It is a nice exercise to prove that $f^{*}(E)\to X$ is really a real vector bundle, of the same dimension as $E$, and each $\widetilde{f}:f^{*}(E)_{x}\to E_{f(x)}$ is an isomorphism on fibres. If $X$ is a subset of $Y$ then $f^{*}(E)$ is called the restriction of the bundle to $X$ and sometimes it is written $E|_{X}$.
@@
On the other hand, the tangent bundle of $S^{2}$ has the property that every section $s:S^{2}\to T(S^{2})$ must be zero at some point. [^62] If this were not the case then you could comb the hair of a coconut.
@@
Suppose that $E$ and $F$ are real vector bundles over the compact space $X$. We define their sum to be the real vector bundle $E\oplus F$ where the fibre over the point $x$ is $E_{x}\oplus F_{x}$. Similarly, the tensor product of the bundles $E\otimes F$ is the real vector bundle where the fibre over the point $x$ is $E_{x}\otimes_{{\mathbb{R}}}F_{x}$. [^63] This requires more careful explanation but is identical to the complex situation, so please refer to the complex authorities. For example, if $E$ is a real vector bundle over $X$ and $\theta_{1}$ is the one-dimensional trivial bundle, then $E\otimes\theta_{1}\cong E$, generalizing the fact that ${\mathbb{R}}^{n}\otimes_{\mathbb{R}}{\mathbb{R}}\cong{\mathbb{R}}^{n}$.
Finally let $Vect_{k}^{\mathbb{R}}(X)$ denote the set of isomorphism classes of real vector bundles of dimension $k$ and $Vect^{\mathbb{R}}(X)$ denote the union of the $Vect_{k}^{\mathbb{R}}(X)$.
It is easy to see that $Vect^{\mathbb{R}}(X)$ has natural addition and multiplication operations, and these give $Vect^{\mathbb{R}}(X)$ the structure of an abelian semiring with unit $\theta_{1}$. It turns out that we can calculate it as well in homotopy terms.
It is best to restrict at this point to bundles over compact spaces. One very important consequence of this restriction is the following fact: If $E$ is any finite-dimensional real vector bundle over a compact space $X$ then there exists another finite-dimensional real vector bundle $F$ over $X$ such that $E\oplus F$ is a trivial bundle.646464cf. Atiyah [7] p. 27, Cor. 1.4.14.
@@
$X,BO_{k}]^{u}\overset{\cong}{\longrightarrow}Vect_{k}^{\mathbb{R}}(X)$
@@
    $X,BU_{k}]^{u}\cong Vect_{k}^{\mathbb{C}}(X).$
@@
$c:Vect_{\mathbb{R}}^{n}(X)\longrightarrow Vect_{\mathbb{C}}^{n}(X)$

### v3

@@ 6. The Homogeneous Spaces $Sp_{n}/(Sp_{k}\times Sp_{n-k})$
We will see in the next section that $\pi_{j}(Sp_{n})=0$ for $j\leq 2$. This implies that in Sequence (12.1) two out of every three groups is zero, and hence
@@ 8. The Homogeneous Spaces $Sp_{n}/U_{n}$ and $Sp_{n}/SU_{n}$
The spaces $Sp_{n}/U_{n}$ are path-connected. Sequence (12.2) gives an exact sequence
@@
For $Sp_{n}/SU_{n}$ Sequence (12.2) gives an exact sequence
@@ 9. The Homogeneous Spaces $U_{n}/O_{n}$, $SU_{n}/SO_{n}$ and $U_{n}/SO_{n}$
First consider $U_{n}/O_{n}$. As $O_{n}$ is not path-connected we use Sequence (12.1) and obtain
@@
shows that the map $\alpha:\pi_{1}(O_{2})\to\pi_{1}(U_{2})$ is an isomorphism. Thus Sequence (12.1) reduces down to
@@
$\pi_{1}(U_{n}/O_{n})\cong{\mathbb{Z}}_{2}\qquad\text{and}\qquad\pi_{2}(U_{n}/O_{n})=0.$
@@
The map $\alpha=0$ and so it follows that $\pi_{2}(U_{n}/O_{n})={\mathbb{Z}}_{2}$ and that $\pi_{1}(U_{n}/O_{n})$ is either isomorphic to $\mathbb{Z}$} or to ${\mathbb{Z}}\oplus{\mathbb{Z}}_{2}$. We will settle this point in a moment.
@@
Turning to Sequence (12.2) we obtain
@@
Finally, consider $U_{n}/SO_{n}$. The space $U_{n}$ is path-connected. If $n=2$ then Sequence (12.2) gives us
@@
$\pi_{1}(U_{n}/SO_{n})\cong{\mathbb{Z}}\qquad\text{and}\qquad\pi_{2}(U_{n}/SO_{n})\cong{\mathbb{Z}}_{2}$
@@
The map $\gamma$ is surjective and hence $\pi_{0}(O_{n}/(O_{k}\times O_{n-k}))=0$. A similar and easier argument works for $SO$ and hence
$\pi_{0}(O_{n}/(O_{k}\times O_{n-k}))=0\qquad\text{and}\qquad\pi_{0}(SO_{n}/(SO_{k}\times SO_{n-k}))=0$
@@
    $\pi_{1}(O_{2}/(O_{1}\times O_{1}))\cong{\mathbb{Z}}\qquad\text{and}\qquad\pi_{2}(O_{2}/(O_{1}\times O_{1})=0.$
@@
    $\pi_{1}(SO_{2}/(SO_{1}\times SO_{1}))\cong{\mathbb{Z}}\qquad\text{and}\qquad\pi_{2}(SO_{2}/(SO_{1}\times SO_{1})=0.$
@@
    with $p$ surjective, and this implies that
    $\pi_{1}(RP^{3}/S^{1})=0\qquad\text{and}\qquad\pi_{2}(RP^{3}/S^{1})={\mathbb{Z}}$
@@
    $\pi_{1}(O_{3}/O_{2})=0\qquad\text{and}\qquad\pi_{2}(O_{3}/O_{2})={\mathbb{Z}}$
@@
    $\pi_{1}(SO_{3}/SO_{2})=0\qquad\text{and}\qquad\pi_{2}(SO_{3}/SO_{2})={\mathbb{Z}}$
@@
    $\pi_{1}(O_{3}/(O_{1}\times O_{2})={\mathbb{Z}}_{2}\qquad\text{and}\qquad\pi_{2}(O_{3}/(O_{1}\times O_{2})\cong{\mathbb{Z}}.$
@@
    $\pi_{1}(SO_{3}/(SO_{1}\times SO_{2})=0\qquad\text{and}\qquad\pi_{2}(SO_{3}/(SO_{1}\times SO_{2})\cong{\mathbb{Z}}.$
@@
    Applying Sequence (12.1) to $O_{4}/(O_{2}\times O_{2})$ gives us
@@
    $0\to\pi_{2}(O_{4}/(O_{2}\times O_{2}))\to{\mathbb{Z}}^{2}\twoheadrightarrow{\mathbb{Z}}_{2}\overset{p}{\longrightarrow}\pi_{1}(O_{4}/(O_{2}\times O_{2}))\to{\mathbb{Z}}_{2}\oplus{\mathbb{Z}}_{2}\to{\mathbb{Z}}_{2}\to 0$
@@
    $0\to\pi_{2}(O_{4}/(O_{2}\times O_{2}))\to{\mathbb{Z}}^{2}\twoheadrightarrow{\mathbb{Z}}_{2}\to 0$
@@
    $\pi_{1}(O_{4}/(O_{2}\times O_{2}))\cong{\mathbb{Z}}_{2}\qquad\text{and}\qquad\pi_{2}(O_{4}/(O_{2}\times O_{2}))\cong{\mathbb{Z}}^{2}$
@@
    $0\to\pi_{2}(SO_{4}/(SO_{2}\times SO_{2}))\longrightarrow\pi_{1}(SO_{2}\times SO_{2})\twoheadrightarrow\pi_{1}(SO_{4})\longrightarrow\pi_{1}(SO_{4}/(SO_{2}\times SO_{2}))\to 0$
@@
    $0\to\pi_{2}(SO_{4}/(SO_{2}\times SO_{2}))\longrightarrow{\mathbb{Z}}^{2}\twoheadrightarrow{\mathbb{Z}}_{2}\overset{p}{\longrightarrow}\pi_{1}(SO_{4}/(SO_{2}\times SO_{2}))\to 0$
@@
    $\pi_{1}(SO_{4}/(SO_{2}\times SO_{2}))\cong 0\qquad\text{and}\qquad\pi_{2}(SO_{4}/(SO_{2}\times SO_{2}))\cong{\mathbb{Z}}^{2}$
@@
    The map $\bar{\gamma}$ is surjective, and hence $\zeta$ is the zero map. The map $\gamma$ is also surjective. Thus
    $\pi_{1}(O_{n}/(O_{k}\times O_{n-k}))\cong{\mathbb{Z}}_{2}\qquad\text{and}\qquad\pi_{2}(O_{n}/(O_{k}\times O_{n-k}))\cong{\mathbb{Z}}_{2}.$
@@
    The case of $SO_{n}/(SO_{k}\times SO_{n-k})$ is much simpler. The space $SO(n)$ is path-connected so Sequence (12.2) gives us
@@
    $0\to\pi_{2}(SO_{n}/(SO_{k}\times SO_{n-k}))\longrightarrow{\mathbb{Z}}_{2}\oplus{\mathbb{Z}}_{2}\twoheadrightarrow{\mathbb{Z}}_{2}\to\pi_{1}(SO_{n}/(SO_{k}\times SO_{n-k}))\to 0.$
@@
    $\pi_{1}(SO_{n}/(SO_{k}\times SO_{n-k}))=0\qquad\text{and}\qquad\pi_{2}(SO_{n}/(SO_{k}\times SO_{n-k}))={\mathbb{Z}}_{2}.$
@@ 13. The Basics: Real Vector Bundles
It is a nice exercise to prove that $f^{*}(E)\to X$ is really a real vector bundle, of the same dimension as $E$, and each $\widetilde{f}:f^{*}(E)_{x}\to E_{f(x)}$ is an isomorphism on fibres. If $X$ is a subset of $Y$ and $f:X\hookrightarrow Y$ is the inclusion map, then $f^{*}(E)$ is called the restriction of the bundle to $X$ and sometimes it is written $E|_{X}$.
@@
On the other hand, the tangent bundle of $S^{2}$ has the property that every section $s:S^{2}\to T(S^{2})$ must be zero at some point. [^63] If this were not the case then you could comb the hair of a coconut. This was establshed by H. Poincaré in 1885.
@@
Suppose that $E$ and $F$ are real vector bundles over the compact space $X$. We define their sum to be the real vector bundle $E\oplus F$ where the fibre over the point $x$ is $E_{x}\oplus F_{x}$. Similarly, the tensor product of the bundles $E\otimes F$ is the real vector bundle where the fibre over the point $x$ is $E_{x}\otimes_{{\mathbb{R}}}F_{x}$. [^64] This requires more careful explanation but is identical to the complex situation, so please refer to the complex authorities. For example, if $E$ is a real vector bundle over $X$ and $\theta_{1}$ is the one-dimensional trivial bundle, then $E\otimes\theta_{1}\cong E$, generalizing the fact that ${\mathbb{R}}^{n}\otimes_{\mathbb{R}}{\mathbb{R}}\cong{\mathbb{R}}^{n}$. Finally let $Vect_{k}^{\mathbb{R}}(X)$ denote the set of isomorphism classes of real vector bundles of dimension $k$ and $Vect^{\mathbb{R}}(X)$ denote the union of the $Vect_{k}^{\mathbb{R}}(X)$.
It is easy to see that the operations $E\oplus F$ and $E\otimes F$ give $Vect^{\mathbb{R}}(X)$ the structure of an abelian semiring with unit $\theta_{1}$. It turns out that we can calculate it as well in homotopy terms.
It is best to restrict at this point to bundles over compact spaces. One very important consequence of this restriction is the following fact: If $E$ is any finite-dimensional real vector bundle over a compact space $X$ then there exists another finite-dimensional real vector bundle $F$ over $X$ such that $E\oplus F$ is a trivial bundle.656565cf. Atiyah [7] p. 27, Cor. 1.4.14.
@@
$X,BO_{k}]^{ub}\overset{\cong}{\longrightarrow}Vect_{k}^{\mathbb{R}}(X)$
@@
    $X,BU_{k}]^{ub}\cong Vect_{k}^{\mathbb{C}}(X).$
@@
$c:Vect^{\mathbb{R}}_{n}(X)\longrightarrow Vect^{\mathbb{C}}_{n}(X)$